OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

distribute -3 to the numbers within the parentheses so -3x - 9 = -3x - 3 - 5 then just add like terms, isolate the variable. :)

OpenStudy (yanasidlinskiy):

@xspoon Do it in the order that I have showed you right before with a different problem.

OpenStudy (yanasidlinskiy):

Show us what you got:) Jadeybells or I will be glad to help you if you et stuck!:) Don't hesitate!!:) Good Luck!!:)

OpenStudy (anonymous):

Well both of the x's cancel out so now the problem looks like this 9 = 3 - 5. What do I do next?

OpenStudy (zzr0ck3r):

since \(9\ne -2\) there is no solutions

OpenStudy (anonymous):

exactly

OpenStudy (zzr0ck3r):

or since \(-9\ne -3-5=-8\) there is no solution...

OpenStudy (anonymous):

Oh okay. Thanks everyone ! (:

OpenStudy (yanasidlinskiy):

Yes. But you forgot the -3. And yes! I agree zzr0ck3r!!:)

OpenStudy (yanasidlinskiy):

No problem!!:)